### v1

# The rigidity of eigenfunctions’ gradient estimates

Abstract.
On compact Riemannian manifolds with non-negative Ricci curvature and smooth (possibly empty), convex (or mean convex) boundary, if the sharp Li-Yau type gradient estimate of an Neumann (or Dirichlet) eigenfunction holds at some non-critical points of the eigenfunction; we show that the manifold is isometric to the product of one lower dimension manifold and a round circle (or a line segment).
Mathematics Subject Classification: 35J15, 58J05, 58J50.
G.Xu was partially supported by NSFC 12141103.
[…]
1. Introduction
For manifolds with non-negative Ricci curvature ($Rc\geq 0$), Cheeger-Gromoll’s splitting theorem [CG] says: the existence of geodesic line results in the isometric splitting of the manifold. It can be shown that the Busemann function with respect to the geodesic line is a harmonic function, whose gradient has constant norm. Then the isometry is induced by the gradient curve of this harmonic function.
Remark 1.1.
Note the square norm of harmonic functions’ gradient is a subharmonic function in manifolds with $Rc\geq 0$. If the maximum norm of harmonic functions’ gradient is obtained at some point $p\in M^{n}$, then the strong maximum principle implies that the norm of harmonic functions’ gradient is constant, which guarantees that the manifold isometrically splits.
Therefore from the pure analytic view point, the above isometric splitting result can be viewed as the rigidity part of harmonic functions’ gradient estimate.
Figure 1. The Figure for Cheeger-Gromoll’s splitting theorem with $M^{n}=N^{n-1}\times\mathbb{R}$
For the global positive Green’s function, the sharp estimate and the related rigidity are known to Colding [Colding]. For the global positive Green’s function $G$ on a complete Riemannian manifold with nonnegative Ricci curvature, it is shown that $|\nabla(G^{\frac{1}{2-n}})|\leq 1$, and $M^{n}$ is isometric to $\mathbb{R}^{n}$ if and only if the equality holds at some point in $M^{n}$.
The harmonic functions in Cheeger-Gromoll’s splitting theorem and Colding’s rigidity are global. The study of local harmonic functions in manifolds with non-negative Ricci curvature, was initiated by Yau’s Liouville theorem [Yau] and Cheng-Yau’s gradient estimate for positive harmonic functions [CY], where the gradient of the logarithm of positive harmonic functions is considered.
In [Xu-growth], the sharp Cheng-Yau type gradient estimate in two dimensional manifolds with $Rc\geq 0$ is obtained. Furthermore, the corresponding rigidity is established in [HXY] as follows: the realization of maximum norm of positive harmonic function’s gradient in geodesic disk implies, the geodesic disk is isometric to an Euclidean disk. However, the sharp Cheng-Yau type gradient estimate is unknown for higher dimension cases so far.
Li [Li] and Li-Yau [LY] applied the gradient estimates of eigenfunctions to study the bound of the eigenvalues on manifolds. The simplest model for eigenfunctions are sine or cosine functions. The Li-Yau type gradient estimate of eigenfunctions (the estimate of $|D\sin^{-1}u|$, where $u$ is one normalized eigenfunction of compact manifold) is firstly obtained in [LY, Theorem $10$].
Through a delicate refined gradient estimate of eigenfunctions, Zhong-Yang [ZY] firstly obtain the sharp lower bound of the first eigenvalue for compact Riemannian manifolds with $Rc\geq 0$. Hang-Wang [HW] answered Sakai’s question in [Sakai], they showed that if the sharp lower bound of the $1$st eigenvalue is obtained, then manifold is isometric to a round circle. Zhong-Yang’s type gradient estimate of eigenfunction is closely related to the solution of suitable ODE (also see [Kroger] and [AX]). There is also log-concavity property of the first eigenfunctions established by two-points method in [AC] (also see [Ni]). For Riemannian manifold satisfying $Rc\geq-(n-1)$, the sharp gradient estimate of the $1$st eigenfunction and the rigidity of sharp $1$st eigenvalue is studied in [SW] (also see [Li12]).
Although Li-Yau type gradient estimate of eigenfunctions is not sharp enough to get the sharp lower bound of the first eigenvalue, the equality in the estimate can be achieved for model spaces; which splits a round circle factor isometrically. In this paper, we study the rigidity part with respect to Li-Yau type gradient estimates of eigenfunctions, which is partially motivated by Remark 1.1.
The main result of this paper is as follows, which can be viewed as an isometric splitting theorem for compact manifolds comparing Cheeger-Gromoll’s splitting Theorem for complete non-compact manifolds.
Theorem 1.2.
For compact Riemannian manifold $(M^{n},g)$ with $Rc\geq 0$ and smooth (possibly empty) boundary . Assume $\Delta u=-\lambda u,\displaystyle\max_{x\in M^{n}}|u(x)|=1$, where $\lambda>0$. If $|\nabla\sin^{-1}u|^{2}(p)=\lambda$ holds at some point $p$, where $|u(p)|<1$. Define $\mathcal{N}(u)=\#(M^{n}-u^{-1}(0))$ as the number of connected components of $M^{n}-u^{-1}(0)$ (the number of nodal domains of $u$); then:
  1. (1)
    . If $\partial M^{n}=\emptyset$, then $M^{n}$ is isometric to $\mathbb{S}^{1}(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}})\times N^{n-1}$, where $\mathbb{S}^{1}(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}})$ is a round circle with radius $\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}$ and $\mathcal{N}(u)=2k$ for some $k\in\mathbb{Z}^{+}$;
    Figure 2. The Figure for Case (1) of Theorem 1.2
  2. (2)
    . If $\partial M^{n}\neq\emptyset$ is convex and $\frac{\partial u}{\partial\vec{n}}\big{|}_{\partial M^{n}}=0$, then $M^{n}$ is isometric to $[0,\frac{\mathcal{N}(u)-1}{\sqrt{\lambda}}\cdot\pi]\times N^{n-1}$ and $\mathcal{N}(u)\geq 2$;
    Figure 3. The Figure for Case (2) of Theorem 1.2
  3. (3)
    . If $\partial M^{n}\neq\emptyset$ is mean convex and $u\big{|}_{\partial M^{n}}=0$; then $M^{n}$ is isometric to $[0,\frac{\mathcal{N}(u)}{\sqrt{\lambda}}\cdot\pi]\times N^{n-1}$ and $\mathcal{N}(u)\geq 1$;
    Figure 4. The Figure for Case (3) of Theorem 1.2
where $N^{n-1}$ is a compact Riemannian manifold with $Rc\geq 0$ and $\partial N^{n-1}=\emptyset$.
Remark 1.3.
The gradient curve of Busemann function is used in [CG72] to show the homeomorphism type of the manifolds. Furthermore the gradient curve of harmonic functions is used to obtain the isometric splitting in [CG]. The proof of the above theorem relies on the gradient curve of $\sin^{-1}u$. However, because of the possible boundary of manifolds and the critical points of eigenfunctions, we firstly obtain a locally splitting result, then extend this splitting to the global manifold.
The organization of this paper is as follows. In Section 2, we firstly give an alternative proof of the sharp gradient estimate of $\sin^{-1}u$, where $u$ is one normalized eigenfunction. The sharp gradient estimate is known in equivalent form (see [Li12], [Yang]), however our method is uniform and consistent in dealing with the corresponding rigidity part later.
In Section 3, we obtain the vanishing Hessian of $\sin^{-1}u$, then using the gradient curve of this function to show the locally splitting in the connected component without critical points of eigenfunctions.
Finally we show the locally splitting can be extended to the globally splitting in Section 4. Note the canonical strong maximum principle (see[GT], [PW]) requires the uniformly elliptic property of the differential operators. Because the elliptic operator related to $\sin^{-1}u$ is degenerate at the critical points of $u$, we can not apply the canonical strong maximum principle directly.
The key observation is that using the Taylor expansion of the Riemannian metric and eigenfunction at the critical points, we find a suitable make-up function, which plays the similar role of the make-up function used in the proof of classical Hopf’s Strong Maximum Principle (see [GT, Lemma $3.4$]). Therefore, we obtain the implication of Strong Maximum Principle for this degenerate elliptic operator, which guarantees the extension of locally splitting.
Figure 5. The Figure of Theorem 4.1 from local splitting to global splitting
2. The sharp gradient estimates for eigenfunctions
We firstly study the behavior of eigenfunctions’ gradient at the boundary, where Neumann boundary condition and smooth convex boundary (respectively Dirichlet boundary condition and smooth, mean convex boundary) are used.
Lemma 2.1.
For compact Riemannian manifold $(M^{n},g)$, assume $\Delta u=-\lambda u$ with $|u|\leq 1$, where $\lambda>0$. If $\partial M^{n}$ is smoothly convex and $\frac{\partial u}{\partial\vec{n}}\big{|}_{\partial M^{n}}=0$, or if $\partial M^{n}$ is smooth, mean convex and $u\big{|}_{\partial M^{n}}=0$; then
$\frac{\operatorname{\partial}|\nabla\sin^{-1}u|^{2}}{\operatorname{\partial}\vec{n}}(x)\leq 0,\quad\quad\quad\quad\forall x\in\partial M^{n}\ \text{satisfying}\ |u(x)|<1;$
where $\vec{n}(x)$ is the outward unit normal vector on $\operatorname{\partial}M^{n}$.
Proof.
Let $\{e_{1},\cdots,e_{n-1},e_{n}\}$ be a local orthogonal frame such that $\{e_{1},\cdots,e_{n-1}\}$ are tangent to $\operatorname{\partial}M^{n}$ and $e_{n}=\vec{n}$ is the outward normal vector.
In the rest argument, we assume $x\in\partial M^{n}$ satisfying $|u(x)|<1$. If $\partial M^{n}$ is smooth, convex and $\frac{\partial u}{\partial\vec{n}}\big{|}_{\partial M^{n}}=0$, then at $x$,
$\nabla_{\vec{n}}|\nabla u|^{2}$ $=\sum_{i=1}^{n}2\nabla_{\vec{n}}\nabla_{i}u\cdot\nabla_{i}u=\sum_{i=1}^{n}-2(\nabla_{i}\vec{n})u\cdot\nabla_{i}u,$
$=\sum_{1\leq i,j\leq n}-2\langle\nabla_{i}\vec{n},j\rangle\nabla_{j}u\cdot\nabla_{i}u=-2\operatorname{\langle}\nabla_{\nabla u}\vec{n},\nabla u\operatorname{\rangle}\leq 0.$
by the convexity of $\operatorname{\partial}M^{n}$.
If $\partial M^{n}$ is smooth, mean convex and $u\big{|}_{\partial M^{n}}=0$, then note $\Delta u(x)=-\lambda u(x)=0$, we have
$\begin{split}\nabla_{\vec{n}}|\nabla u|^{2}&=\sum_{i=1}^{n}2\nabla_{\vec{n}}\nabla_{i}u\cdot\nabla_{i}u=2\nabla_{\vec{n}}\nabla_{\vec{n}}u\cdot\nabla_{\vec{n}}u=2\nabla^{2}u(e_{n},e_{n})\cdot\nabla_{\vec{n}}u\\ &=2(\Delta u-\sum_{i=1}^{n-1}\nabla^{2}u(e_{i},e_{i}))\cdot\nabla_{\vec{n}}u=-2\sum_{i=1}^{n-1}\nabla^{2}u(e_{i},e_{i})\cdot\nabla_{\vec{n}}u\\ &=-2\sum_{i=1}^{n-1}(\nabla_{i}\nabla_{i}u-(\nabla_{e_{i}}e_{i})u)\cdot\nabla_{\vec{n}}u=2\sum_{i=1}^{n-1}\operatorname{\langle}\nabla_{e_{i}}e_{i},\nabla u\operatorname{\rangle}\cdot\nabla_{\vec{n}}u\\ &=2\sum_{i=1}^{n-1}\operatorname{\langle}\nabla_{e_{i}}e_{i},e_{n}\operatorname{\rangle}\cdot(\nabla_{\vec{n}}u)^{2}\leq 0,\end{split}$
where the last inequality comes from the mean-convex property of $\operatorname{\partial}M^{n}$.
Now we have
$\nabla_{\vec{n}}|\nabla\sin^{-1}u|^{2}(x)=\frac{(\nabla_{\vec{n}}|\nabla u|^{2})(1-u^{2})-|\nabla u|^{2}(-2u\nabla_{\vec{n}}u)}{(1-u^{2})^{2}}=\frac{\nabla_{\vec{n}}|\nabla u|^{2}}{1-u^{2}}\leq 0.$
Notation 2.2.
For $u\in C^{\infty}(M^{n})$, for any domain $\Omega\subseteq M^{n}$ with $|\nabla u|(x)\neq 0,|u(x)|<1$ in $\Omega$; we define $\mathscr{L}_{u}:C^{\infty}(\Omega)\rightarrow C^{\infty}(\Omega)$ as follows:
$\mathscr{L}_{u}(f)(x)\vcentcolon=\Delta f(x)+\nabla f(x)\cdot\Psi(x)-\frac{2|\nabla u|^{2}}{(1-u^{2})^{2}}f(x),\quad\quad\forall f\in C^{\infty}(\Omega);$
(2.1) $\Psi(x)=\frac{\nabla u}{1-u^{2}}\cdot\{-\frac{1}{2}\frac{(1-u^{2})^{2}}{|\nabla u|^{4}}\cdot\nabla u\cdot\nabla\Big{(}\frac{|\nabla u|^{2}}{1-u^{2}}(x)\Big{)}-2u\}.$
The following lemma is the key differential inequality for make-up function $F(x)=\frac{|\nabla u|^{2}}{1-u^{2}}(x)-\lambda$, which will be crucially used in the proof Theorem 4.1.
Lemma 2.3.
Suppose $M^{n}$ is a compact Riemannian manifold with $Rc\geq 0$, let $u$ be an eigenfunction with $\Delta u=-\lambda u$ in $M^{n}$, where $\lambda>0$. For any $x\in M^{n}$ with $|u(x)|<1$ and $|\nabla u(x)|\neq 0$, let $F(x)=\frac{|\nabla u|^{2}}{1-u^{2}}(x)-\lambda$, we have
$\mathscr{L}_{u}(F)(x)\geq 0,$
Proof: Let $\{e_{1},\cdots,e_{n}\}$ be an orthonormal basis at $T_{x}M$ with $\nabla_{e_{i}}e_{j}(x)=0$, $1\leq i,j\leq n$, where $e_{1}=\frac{\nabla u}{|\nabla u|}(x)$. Denote $u_{i}(x)=\nabla_{e_{i}}u(x)$ and $u_{ij}=\nabla_{e_{j}}\nabla_{e_{i}}u(x)$, then $u_{1}(x)=|\nabla u|(x)$ and $u_{i}(x)=0$, where $2\leq i\leq n$.
Direct computation yields
$\nabla F=2|\nabla u|\cdot\Big{(}\frac{\nabla^{2}u\cdot e_{1}}{1-u^{2}}+\frac{|\nabla u|^{2}\cdot u\cdot e_{1}}{(1-u^{2})^{2}}\Big{)}.$
Then we have
$u_{11}(x)=\frac{(1-u^{2})\cdot\nabla F\cdot e_{1}}{2|\nabla u|}-\frac{|\nabla u|^{2}\cdot u}{(1-u^{2})}.$
And we get
(2.2) $|\nabla^{2}u|^{2}\geq u_{11}^{2}=\frac{|\nabla u|^{4}\cdot u^{2}}{(1-u^{2})^{2}}+\frac{(1-u^{2})\nabla F\cdot e_{1}}{2|\nabla u|}\cdot\Big{\{}\frac{(1-u^{2})\nabla F\cdot\nabla u}{2|\nabla u|^{2}}-2\frac{|\nabla u|^{2}u}{1-u^{2}}\Big{\}}.$
From the Bochner’s formula and (2.2), we can get
$\Delta F$ $\geq-2\nabla F\cdot\nabla[\ln(1-u^{2})]+\frac{2\cdot\Big{\{}|\nabla u|^{2}(|\nabla u|^{2}-\lambda)+(1-u^{2})\cdot(|\nabla^{2}u|^{2}-\lambda|\nabla u|^{2})\Big{\}}}{(1-u^{2})^{2}}\cdot$
$\geq\frac{2|\nabla u|^{2}}{(1-u^{2})^{2}}F-\nabla F\cdot\Psi(x);$
where $\Psi$ is defined as in (2.1). Then the conclusion follows. ∎
Now we are ready to show the sharp gradient estimate for (Neumann or Dirichlet) eigenfunctions in a uniform way.
Theorem 2.4.
For compact Riemannian manifold $(M^{n},g)$ with $Rc\geq 0$, assume $\Delta u=-\lambda u,\displaystyle\max_{x\in M^{n}}|u(x)|=1$, where $\lambda>0$. If $\partial M^{n}$ is smooth, convex and $\frac{\partial u}{\partial\vec{n}}\big{|}_{\partial M^{n}}=0$, or if $\partial M^{n}$ is smooth, mean convex and $u\big{|}_{\partial M^{n}}=0$; then
$|\nabla u|^{2}+\lambda u^{2}\leq\lambda.$
Remark 2.5.
The Neumann case is originally proved in [LY], and the Dirichlet case is proved in [Yang, Theorem $3.2$].
Proof.
Step (1). Let $u_{\operatorname{\epsilon}}=\frac{u}{1+\operatorname{\epsilon}}$, where $\operatorname{\epsilon}>0$, then $|u_{\operatorname{\epsilon}}|<1$. We define $F_{\operatorname{\epsilon}}(x)=\frac{|\nabla u_{\operatorname{\epsilon}}|^{2}}{1-u_{\operatorname{\epsilon}}^{2}}-\lambda$. Suppose $F_{\operatorname{\epsilon}}(x_{0})=\max\limits_{M^{n}}\frac{|\nabla u_{\operatorname{\epsilon}}|^{2}}{1-u_{\operatorname{\epsilon}}^{2}}$. There are two cases :
If $x_{0}\in M^{n}\backslash\operatorname{\partial}M^{n}$, then we get $\nabla F_{\operatorname{\epsilon}}(x_{0})=0$.
If $x_{0}\in\operatorname{\partial}M^{n}$, we can choose an orthonormal basis $\{e_{1},\cdots,e_{n}\}$ in the neighborhood $U$ of $x_{0}$ such that $e_{n}(x)=\vec{n}(x)$ for all $x\in\operatorname{\partial}M^{n}\cap U$ and $\nabla_{e_{n}}e_{i}(x)=0$ for all $x\in U$, $1\leq i\leq n$.
From the definition of $x_{0}$, we know
(2.3) $\nabla_{i}F_{\operatorname{\epsilon}}(x_{0})=0,\,i=1,\cdots,n-1\,\,\text{and}\,\nabla_{\vec{n}}F_{\operatorname{\epsilon}}(x_{0})\geq 0.$
Note $\nabla_{\vec{n}}F_{\operatorname{\epsilon}}=\nabla_{\vec{n}}|\nabla\sin^{-1}u_{\epsilon}|^{2}$, from Lemma 2.1, we get that $\nabla_{\vec{n}}F_{\operatorname{\epsilon}}(x_{0})\leq 0$. Now $\nabla_{\vec{n}}F_{\operatorname{\epsilon}}(x_{0})=0$ by (2.3). Therefore, $\nabla F_{\operatorname{\epsilon}}(x_{0})=0$ for $x_{0}\in\operatorname{\partial}M^{n}$.
Then we obtain
(2.4) $\nabla F_{\operatorname{\epsilon}}(x_{0})=0.$
Step (2). Then from the choice of $x_{0}$, for any $v\in ST_{x_{0}}M^{n}$ and $\gamma_{x_{0}}(tv)=\exp_{x_{0}}(tv)$, we get
$\nabla^{2}_{v,v}F_{\operatorname{\epsilon}}(x_{0})$ $=2\lim_{t\rightarrow 0}\frac{F_{\operatorname{\epsilon}}(\gamma_{x_{0}}(tv))-F_{\operatorname{\epsilon}}(x_{0})-\nabla_{v}F_{\operatorname{\epsilon}}(x_{0})\cdot t}{t^{2}}=2\lim_{t\rightarrow 0}\frac{F_{\operatorname{\epsilon}}(\gamma_{x_{0}}(tv))-F_{\operatorname{\epsilon}}(x_{0})}{t^{2}}$
(2.5) $\leq 0.$
We suppose $\nabla u_{\operatorname{\epsilon}}(x_{0})\neq 0$ in the rest argument, otherwise $|\nabla u_{\operatorname{\epsilon}}|\equiv 0$ on $M^{n}$, and the conclusion is trivial.
Note $u_{\epsilon}$ is an eigenfunction with $|u_{\epsilon}|<1$. From Lemma 2.3, we get
$\mathscr{L}_{u_{\epsilon}}F_{\epsilon}(x_{0})\geq 0.$
From the above, (2.4) and (2.5), we get $F_{\operatorname{\epsilon}}(x_{0})\leq 0$, which implies $|\nabla u_{\operatorname{\epsilon}}|^{2}(x_{0})+\lambda u_{\operatorname{\epsilon}}^{2}(x_{0})\leq\lambda$. By the definition of $x_{0}$, we know $|\nabla u_{\operatorname{\epsilon}}|^{2}+\lambda u_{\operatorname{\epsilon}}^{2}(x)\leq\lambda$ for all $x\in M^{n}$. Let $\operatorname{\epsilon}\to 0$, we get $|\nabla u|^{2}+\lambda u^{2}\leq\lambda$. ∎
3. The local splitting induced by eigenfunctions
Definition 3.1.
For $u\in C^{\infty}(M^{n})$ and $p\in M^{n}$, if $\nabla u(p)\neq 0$, let $M_{u,p}$ be the connected component of $M^{n}\backslash\mathscr{C}(u)$ containing $p$, where $\mathscr{C}(u)=\{x\in M^{n}:|\nabla u|(x)=0\}$. When the context is clear, we also use $M_{p}$ for simplicity.
Proposition 3.2.
For compact Riemannian manifold $(M^{n},g)$ with $Rc\geq 0$. Assume $\Delta u=-\lambda u,\displaystyle\max_{x\in M^{n}}|u(x)|=1$, where $\lambda>0$. If $|\nabla\sin^{-1}u|^{2}(p)=\lambda$ holds at some point $p$ with $|u(p)|<1$ , then $N^{n-1}\vcentcolon=u^{-1}(u(p))\cap M_{p}$ is connected and $\partial N^{n-1}=\emptyset$. Furthermore,
  1. (N)
    . If $\partial M^{n}=\emptyset$, then $M_{p}$ is isometric to $N^{n-1}\times\left(-\frac{\pi}{2\sqrt{\lambda}},\frac{\pi}{2\sqrt{\lambda}}\right)$ and $\partial N^{n-1}=\emptyset$.
  2. (NB)
    . If $\partial M^{n}\neq\emptyset$ is smooth, convex and $\frac{\partial u}{\partial\vec{n}}\big{|}_{\partial M^{n}}=0$, then $M_{p}$ is isometric to $N^{n-1}\times\left(-\frac{\pi}{2\sqrt{\lambda}},\frac{\pi}{2\sqrt{\lambda}}\right)$.
  3. (DB)
    . If $\partial M^{n}\neq\emptyset$ is smooth, mean convex and $u\big{|}_{\partial M^{n}}=0$; then $\partial N^{n-1}=\emptyset$. And
    1. (DB-a)
      If $M_{p}\cap\partial M^{n}=\emptyset$, then $M_{p}$ is isometric to $N^{n-1}\times\left(-\frac{\pi}{2\sqrt{\lambda}},\frac{\pi}{2\sqrt{\lambda}}\right)$.
    2. (DB-b)
      If $M_{p}\cap\partial M^{n}\neq\emptyset$, then $M_{p}$ is isometric to $N^{n-1}\times\left(-\frac{\pi}{2\sqrt{\lambda}},0\right]$.
Proof: Step (1). First note that for any $x\in M_{p}$, we have $|\nabla u|(x)\neq 0$ and $|u|(x)<1$ by the definition of $M_{p}$ and Theorem 2.4.
Let $V=\{x\in M_{p}\,\,|\,\,|\nabla\sin^{-1}u|(x)=\sqrt{\lambda}\}$, which is nonempty since $p\in V$. It’s evident that $V$ is a closed set in $M_{p}$.
By Lemma 2.3, we have
$\mathscr{L}_{u}(F)(x)\geq 0,\quad\quad\quad\forall x\in M_{p};$
where $F(x)=|\nabla\sin^{-1}u|^{2}(x)-\lambda$.
Given $x_{1}\in V$, note $F$ attains its maximum at $x_{1}$ with $F(x_{1})=0$ by Theorem 2.4. Note $\frac{\partial F}{\partial\vec{n}}\big{|}_{\partial M^{n}}\leq 0$ by Lemma 2.1, we apply the Strong Maximum Principle (see [GT, Lemma $3.4$ and Theorem $3.5$]) to get $F\big{|}_{B_{x_{1}}(r)\cap M_{p}}\equiv 0$. Therefore $(B_{x_{1}}(r)\cap M_{p})\subset V$, and $V$ is open in $M_{p}$.
Because $M_{p}$ is connected , we conclude that $|\nabla\sin^{-1}u|(x)=\sqrt{\lambda}$ for any $x\in M_{p}$.
Let $\Theta(x)=\frac{1}{\sqrt{\lambda}}\sin^{-1}u(x)$ for $x\in M_{p}$, then direct computation combining $|\nabla\Theta|(x)=1,\Delta u=-\lambda u$ yields
$\Delta\Theta=0,\quad\quad\quad\quad\forall x\in M_{p}.$
From the Bochner formula and $Rc\geq 0$, we can infer that
$\begin{split}|\nabla^{2}\Theta|^{2}&=\frac{1}{2}\Delta|\nabla\Theta|^{2}-\langle\nabla\Delta\Theta,\nabla\Theta\rangle-Rc(\nabla\Theta,\nabla\Theta)\leq 0.\end{split}$
Hence $|\nabla^{2}\Theta|\equiv 0$ and $|\nabla\Theta|\equiv 1$ in $M_{p}$.
Step (2). Define $\mathcal{L}_{t}=\big{\{}\ x\in M_{p}:\Theta(x)=t\big{\}}$, from $|\nabla\Theta|\equiv 1$ and the Implicit Function Theorem, we know that $\mathcal{L}_{t}$ is $(n-1)$-dim submanifold of $M^{n}$ if $\mathcal{L}_{t}\neq\emptyset$.
For any $x\in M_{p}$, let $\gamma_{x}$ be the gradient curve of $\nabla\Theta$ with $\gamma_{x}(0)=x$ in $M_{p}$. Define
$T_{x}\vcentcolon=\sup\{t\geq 0:\gamma_{x}(t)\in M_{p}\},\quad\quad\hat{T}_{x}\vcentcolon=\sup\{t\geq 0:\gamma_{x}(-t)\in M_{p}\},\quad\quad\forall x\in M_{p}.$
Now we prove the following claim:
Claim 3.3.
$\mathcal{P}(x)=\gamma_{x}\cap\mathcal{L}_{0}:M_{p}\rightarrow\mathcal{L}_{0}$ is well-defined for any $x\in M_{p}$.
Proof: [Proof of Claim 3.3] If $\Theta(x)=0$, then $\mathcal{P}(x)=x$.
Now we consider the case $\Theta(x)<0$. We firstly show $T_{x}\geq-\Theta(x)$. By contradiction, assume $T_{x}<-\Theta(x)$. Then by $|\nabla\Theta|\leq 1$, we have
$-\frac{\pi}{2\sqrt{\lambda}}<\Theta(x)\leq\Theta(\gamma_{x}(t))\leq\Theta(x)+t\leq\Theta(x)+T_{x}<0,\quad\quad\forall t\in[0,T_{x}).$
Using $|\nabla\Theta|\equiv 1$ in $M_{p}$ and $u(z)=\sin(\sqrt{\lambda}\Theta(z))$ for any $z\in M_{p}$, we get
(3.1) $|\nabla u|(\gamma_{x}(t))\geq\sqrt{\lambda}\cdot\sqrt{1-\sin^{2}(\sqrt{\lambda}\Theta(x))}>0,\quad\forall t\in[0,T_{x}).$
From (3.1), we get
(3.2) $\lim_{t\rightarrow T_{x}}|\nabla u|(\gamma_{x}(t))\geq\sqrt{\lambda}\cdot\sqrt{1-\sin^{2}(\sqrt{\lambda}\Theta(x))}>0,$
which implies $\gamma_{x}(T_{x})$ is well-defined and $\gamma_{x}(T_{x})\in M_{p}$.
Furthermore,
(3.3) $u(\gamma_{x}(t))=\sin(\sqrt{\lambda}\Theta(\gamma_{x}(t)))\leq\sin(\sqrt{\lambda}[\Theta(x)+T_{x}])<0,\quad\quad\forall t\in[0,T_{x}).$
For case (DB), from (3.3), we know $\gamma_{x}(T_{x})\notin\partial M^{n}$. Then there is $\delta>0$ such that $\gamma_{x}(T_{x}+\delta)\in M_{p}$.
For case (NB), from (3.2), we know $|\nabla u|\left(\gamma_{x}(T_{x})\right)>0$.
  1. (a)
    . If $\gamma_{x}(T_{x})\notin\partial M^{n}$, there is $\delta>0$ such that $\gamma_{x}(T_{x}+\delta)\in M_{p}$.
  2. (b)
    . If $\gamma_{x}(T_{x})\in\partial M^{n}$, from $\gamma_{x}^{\prime}(T_{x})=\nabla\Theta\left(\gamma_{x}(T_{x})\right)=\frac{\nabla u}{|\nabla u|}\left(\gamma_{x}(T_{x})\right)\neq 0$ and $\frac{\operatorname{\partial}u}{\operatorname{\partial}\vec{n}}\left(\gamma_{x}(T_{x})\right)=0$; we obtain $\gamma_{x}^{\prime}(T_{x})\in T_{\gamma_{x}(T_{x})}\operatorname{\partial}M^{n}$.
    Combining $\operatorname{\partial}M^{n}$ is smooth and $(\nabla u\cdot\vec{n})\big{|}_{\partial M^{n}}=0$, we can extend $\gamma_{x}$ (which is the gradient curve of $u$) in $\partial M^{n}$. Hence there is $\delta>0$ such that $\gamma_{x}(T_{x}+\delta)\in M_{p}$.
By the above argument, there is $\delta>0$ such that $\gamma_{x}(T_{x}+\delta)\in M_{p}$. It contradicts the definition of $T_{x}$. Therefore $T_{x}\geq-\Theta(x)$.
Next we prove that $\gamma_{x}(-\Theta(x))\in M_{p}$. Like (3.1), we have
(3.4) $|\nabla u|(\gamma_{x}(t))\geq\sqrt{\lambda}\cdot\sqrt{1-\sin^{2}(\sqrt{\lambda}\Theta(x))}>0,\quad\quad\forall t\in[0,-\Theta(x)).$
This implies $\gamma_{x}(-\Theta(x))\in M_{p}$ is well-defined.
Now $\mathcal{P}(x)=\gamma_{x}(-\Theta(x))$ is well defined.
If $\Theta(x)>0$, consider the gradient curve of $-\nabla\Theta$ as above, we can get that $\mathcal{P}(x)$ is also well-defined. By all the above, we know that $\mathcal{P}(x)$ is well-defined for $x\in M_{p}$. ∎
Step (3). Now for cases (N), (NB) and (DB-a), we will show that for each $x\in\mathcal{L}_{0}$, there is
(3.5) $T_{x}=\hat{T}_{x}=\frac{\pi}{2\sqrt{\lambda}}.$
By contradiction, if $T_{x}<\frac{\pi}{2\sqrt{\lambda}}$, then $\Theta(\gamma_{x}(T_{x}))<\frac{\pi}{2\sqrt{\lambda}}$.
We get $|\nabla u|(\gamma_{x}(T_{x}))=\sqrt{\lambda}\cdot\sqrt{1-u^{2}(\gamma_{x}(T_{x}))}>0$. Hence there is $\delta>0$ such that $\gamma_{x}(T_{x}+\delta)\in M_{p}$, which contradicts the definition of $T_{x}$.
Similar argument applies on $\hat{T}_{x}$. Therefore (3.5) is proved.
Similarly, for case (DB-b) we can show
(3.6) $T_{x}\cdot\hat{T}_{x}=0,\quad\text{and}\quad T_{x}+\hat{T}_{x}=\frac{\pi}{2\sqrt{\lambda}},\quad\quad\quad\forall x\in\mathcal{L}_{0}\cap\partial M^{n}.$
For cases (N), (NB) and (DB-a), we define the map
(3.7) $\psi(x)=(\Theta(x),\mathcal{P}(x)):M_{p}\rightarrow(-\frac{\pi}{2\sqrt{\lambda}},\frac{\pi}{2\sqrt{\lambda}})\times\mathcal{L}_{0}.$
Because $x=\gamma_{\mathcal{P}(x)}(\Theta(x))$ and (3.5), we get that $\psi$ is a bijective map. Hence $\psi$ is a diffeomorphism between $M_{p}$ and $(-\frac{\pi}{2\sqrt{\lambda}},\frac{\pi}{2\sqrt{\lambda}})\times\mathcal{L}_{0}$.
For case (DB-b), we get that the map $\psi(x)=(\Theta(x),\mathcal{P}(x))$ is a diffeomorphism between $M_{p}$ and $(-\frac{\pi}{2\sqrt{\lambda}},0]\times\mathcal{L}_{0}$ (or $[0,\frac{\pi}{2\sqrt{\lambda}})\times\mathcal{L}_{0}$).
Step (4). Assume $\xi_{1}^{\prime}(0)=v_{1}\in T_{x}M^{n}$, now using $|\nabla^{2}\Theta|\equiv 0$ and the definition of $\mathcal{P}$, we get
$\langle\psi_{*}v_{1},\psi_{*}v_{1}\rangle$ $=\frac{d}{dt}\psi(\xi_{1})\cdot\frac{d}{dt}\psi(\xi_{1})=\langle\frac{d}{dt}(\Theta(\xi_{1}(t)),\mathcal{P}(\xi_{1}(t))),\frac{d}{dt}(\Theta(\xi_{1}(t)),\mathcal{P}(\xi_{1}(t)))\rangle$
$=\langle\frac{d}{dt}\Theta(\xi_{1}(t)),\frac{d}{dt}\Theta(\xi_{1}(t))\rangle+\langle\frac{d}{dt}\mathcal{P}(\xi_{1}(t)),\frac{d}{dt}\mathcal{P}(\xi_{1}(t))\rangle$
$=\langle\nabla\Theta\cdot v_{1},\nabla\Theta\cdot v_{1}\rangle+\langle\nabla\mathcal{P}\cdot v_{1},\nabla\mathcal{P}\cdot v_{1}\rangle=\langle v_{1},v_{1}\rangle.$
Hence $\psi$ is a local isometry, the conclusion $\psi$ is an isometry follows from all the above.
Therefore $N^{n-1}=u^{-1}(u(p))\cap M_{p}$ is isometric to $\mathcal{L}_{0}$. Because $M_{p}$ is connected, we get $N^{n-1}$ is connected, by the above isometry between $M_{p}$ and the product of $N^{n-1}$ with an interval.
Also we obtain $\partial N^{n-1}=\emptyset$ by the above isometry and $\partial M^{n}$ is smooth. ∎
4. The global splitting
Theorem 4.1.
For compact Riemannian manifold $(M^{n},g)$ with $Rc\geq 0$ and smooth (possibly empty) boundary . Assume $\Delta u=-\lambda u,\displaystyle\max_{x\in M^{n}}|u(x)|=1$, where $\lambda>0$. If $|\nabla\sin^{-1}u|^{2}(p)=\lambda$ holds at some point $p$, where $|u(p)|<1$. Define $\mathcal{N}(u)=\#(M^{n}-u^{-1}(0))$ as the number of connected components of $M^{n}-u^{-1}(0)$ (the number of nodal domains of $u$); then:
  1. (1)
    . If $\partial M^{n}=\emptyset$, then $M^{n}$ is isometric to $\mathbb{S}^{1}(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}})\times N^{n-1}$, where $\mathbb{S}^{1}(\frac{\mathcal{N}(u)}{2\sqrt{\lambda}})$ is a round circle with radius $\frac{\mathcal{N}(u)}{2\sqrt{\lambda}}$ and $\mathcal{N}(u)=2k$ for some $k\in\mathbb{Z}^{+}$;
  2. (2)
    . If $\partial M^{n}\neq\emptyset$ is convex and $\frac{\partial u}{\partial\vec{n}}\big{|}_{\partial M^{n}}=0$, then $M^{n}$ is isometric to $[0,\frac{\mathcal{N}(u)-1}{\sqrt{\lambda}}\cdot\pi]\times N^{n-1}$ and $\mathcal{N}(u)\geq 2$;
  3. (3)
    . If $\partial M^{n}\neq\emptyset$ is mean convex and $u\big{|}_{\partial M^{n}}=0$; then $M^{n}$ is isometric to $[0,\frac{\mathcal{N}(u)}{\sqrt{\lambda}}\cdot\pi]\times N^{n-1}$ and $\mathcal{N}(u)\geq 1$;
where $N^{n-1}$ is a compact Riemannian manifold with $Rc\geq 0$ and $\partial N^{n-1}=\emptyset$.
Proof: Step (1). Define $N^{n-1}=u^{-1}(-1)\cap\overline{M_{p}}$, which is an $(n-1)$-dim hypersurface in $M^{n}$ by Proposition 3.2 (if $u^{-1}(-1)\cap\overline{M_{p}}=\emptyset$, we define $N^{n-1}=u^{-1}(1)\cap\overline{M_{p}}$, similar argument applies); and $U_{\epsilon}(N^{n-1})=\{x\in M^{n}:d(x,N^{n-1})<\epsilon\}$. Then we have a coordinate chart $\{x_{1},x^{\prime}\}$ for $U_{r_{0}}(N^{n-1})$ locally, where $x^{\prime}=(x_{2},\cdots,x_{n})$ is a coordinate chart for $N^{n-1}$ and
$x_{1}\big{|}_{M_{p}\cap U_{r_{0}}(N^{n-1})}=d(x,N^{n-1}),\quad\quad\quad x_{1}\big{|}_{U_{r_{0}}(N^{n-1})-M_{p}}=-d(x,N^{n-1}),$
where $r_{0}>0$ is to be determined later. We define $U_{r_{0}}^{-}(N^{n-1})\vcentcolon=U_{r_{0}}(N^{n-1})\cap\{x\in M^{n}:x_{1}<0\}$.
Consider the metric $g$ in $U_{r_{0}}(N^{n-1})$, then we can write $g=dx_{1}^{2}+\sum_{2\leq i,j\leq n}f_{ij}(x_{1},x^{\prime})dx_{i}dx_{j}$ locally. By Proposition 3.2, there is $a>0$, such that the coefficients $f_{ij}$ satisfy:
[…]
For any $q\in N^{n-1}$, without loss of generality we can assume $q=(0,0,\cdots,0)$. Now we get
$f_{ij}(0,x^{\prime})=\delta_{ij}+O(|x^{\prime}|).$
Denote $G=\mathrm{det}(f_{ij})_{2\leq i,j\leq n}$, we get
$(\log\sqrt{G})_{x_{1}}=\frac{1}{2}(\log G)_{x_{1}}=O(x_{1}).$
We use $u_{i},u_{ij}$ to denote $\frac{\partial u}{\partial x_{i}},\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}$ in the rest argument, similar notation applies for other functions.
Note $u_{11}(q)=\lambda$ because $u\big{|}_{M_{p}}(x_{1},x^{\prime})=\sin(\sqrt{\lambda}x_{1}-\frac{\pi}{2})$ from Proposition 3.2. Also we have
$u_{\alpha}(q)=0,\quad\quad\text{if}\quad\quad\alpha=(\alpha_{1},\cdots,\alpha_{k})\neq(1,\cdots,1),$
$u_{\beta}(q)=0,\quad\quad\text{if}\quad\quad\beta=(1,\cdots,1)\ \text{and there are odds number of $1$},$
Note $N^{n-1}$ is compact, there is a finite coordinate chart covering of $N^{n-1}$. Using the Taylor’s expansion of $u(x_{1},x^{\prime})$ at $q$ about $x_{1}$, where $q$ can be freely chosen from $N^{n-1}$; we get
(4.1) $u(x)=-1+\frac{1}{2}\lambda x_{1}^{2}+O(|x_{1}|^{4}),\quad\quad\quad\forall x\in U_{r_{0}}(N^{n-1}).$
By (4.1), we can choose $r_{0}>0$ such that
$|\nabla u\big{|}_{\overline{U_{r_{0}}(N^{n-1})}-N^{n-1}}|>0,\quad\quad\text{and}\quad\quad|u\big{|}_{\overline{U_{r_{0}}(N^{n-1})}-N^{n-1}}|<1.$
Step (2). Now we define $F:\overline{U_{r_{0}}^{-}(N^{n-1})}\rightarrow\mathbb{R}$ as follows:
$F(x)\vcentcolon=\left\{\begin{array}[]{rl}&\frac{|\nabla u|^{2}}{1-u^{2}}(x)-\lambda\ ,\quad\quad\quad\quad\quad\quad\quad\quad x\in\overline{U_{r_{0}}^{-}(N^{n-1})}-N^{n-1},\\ &0\ ,\quad\quad\quad\quad\quad\quad\quad x\in N^{n-1}.\end{array}\right.$
From the Taylor’s expansion of $u$ at any point of $N^{n-1}$, we get that $F$ is continuous on $\overline{U_{r_{0}}^{-}(N^{n-1})}$.
Then from (4.1), we have
(4.2) $\frac{-2|\nabla u|^{2}}{(1-u^{2})^{2}}$ $=-\frac{2}{x_{1}^{2}}+O(1).$
Direct computation yields
$\Psi(x)$ $=\frac{\nabla u}{1-u^{2}}\cdot\{-\frac{1}{2}\frac{(1-u^{2})^{2}}{|\nabla u|^{4}}\cdot\nabla u\cdot\nabla\Big{(}\frac{|\nabla u|^{2}}{1-u^{2}}(x)\Big{)}-2u\}$
(4.3) $=\Big{(}\frac{2}{x_{1}}+O(x_{1})\Big{)}\cdot\partial x_{1}.$
Consider $v(x)=x_{1}^{2}-x_{1}$, then $v$ satisfies
(4.4) $v\big{|}_{U_{r_{0}}^{-}(N^{n-1})}\geq 0,\quad\quad v\big{|}_{x_{1}=0}=0,\quad\quad\frac{\partial}{\partial x_{1}}v\big{|}_{x_{1}=0}=-1<0.$
Furthermore, using (4.3) and (4.2), we have
$\mathscr{L}_{u}(v)$ $=\Delta v(x)+\nabla v(x)\cdot\Psi(x)-\frac{2|\nabla u|^{2}}{(1-u^{2})^{2}}v(x)$
$=v_{11}+(\log\sqrt{G})_{x_{1}}v_{1}+\sum_{2\leq i,j\leq n}f^{ij}v_{ij}+\sum_{2\leq i,j\leq n}\{\frac{\operatorname{\partial}f^{ij}}{\operatorname{\partial}x_{i}}+(\log\sqrt{G})_{x_{i}}f^{ij}\}v_{j}$
$+v_{1}\cdot(\frac{2}{x_{1}}+O(x_{1}))-\frac{2}{x_{1}^{2}}v+O(1)v$
$=2+O(x_{1})+(2x_{1}-1)\cdot(\frac{2}{x_{1}}+O(x_{1}))-\frac{2}{x_{1}^{2}}\{x_{1}^{2}-x_{1}\}$
$=4+O(x_{1})>0,$
in the last inequality above we use $x_{1}\leq r_{0}<<1$.
Now there is $r_{0}>0$ and $v\in C^{\infty}(U_{r_{0}}(N^{n-1}))$ satisfying
(4.5) $\mathscr{L}_{u}(v)(x)>0,\quad\quad\quad\forall x\in U_{r_{0}}^{-}(N^{n-1}).$
Step (3). For any $q\in N^{n-1}$, let $\vec{n}=\partial x_{1}$ be the outer normal unit vector of $\partial U_{r_{0}}^{-}(N^{n-1})$ at $q$. Using (4.1) we can compute directly that
(4.6) $\frac{\partial F}{\partial\vec{n}}(q)=0.$
We will show that there is $z\in U_{r_{0}}^{-}(N^{n-1})$ such that $F(z)=F(q)=0$. By contradiction, otherwise
(4.7) $F(z)<F(q)=0,\quad\quad\quad\forall z\in U_{r_{0}}^{-}(N^{n-1}).$
From Lemma 2.3, we have
(4.8) $\mathscr{L}_{u}(F)\geq 0.$
Now we consider $F+\epsilon\cdot v$. Because of (4.7), we can choose $\epsilon>0$ such that
(4.9) $(F+\epsilon\cdot v)\big{|}_{x_{1}=-r_{0}}\leq 0.$
Then we have
$(F+\epsilon\cdot v)\big{|}_{\partial U_{r_{0}}^{-}(N^{n-1})}\leq 0.$
By (4.5), we get $\mathscr{L}_{u}(F+\epsilon\cdot v)\big{|}_{U_{r_{0}}^{-}(N^{n-1})}>0$, then from the Maximum principle, we get
$\sup_{x\in U_{r_{0}}^{-}(N^{n-1})}(F+\epsilon\cdot v)(x)\leq\max\{0,\sup_{x\in\partial U_{r_{0}}^{-}(N^{n-1})}(F+\epsilon\cdot v)(x)\}=0.$
Note $(F+\epsilon\cdot v)(q)=0$, then we obtain
$\frac{\partial}{\partial x_{1}}\big{|}_{x=q}(F+\epsilon\cdot v)(x)\geq 0.$
Combining (4.4), this yields
$\frac{\partial}{\partial x_{1}}\big{|}_{x=q}F(x)\geq-\epsilon\cdot\frac{\partial}{\partial x_{1}}\big{|}_{x=q}v(x)=\epsilon>0.$
This contradicts (4.6).
Therefore, we find $z\in U_{r_{0}}^{-}(N^{n-1})$ such that $F(z)=F(q)=0$, which implies $F(z)=\max_{x\in\overline{U_{r_{0}}^{-}(N^{n-1})}}F(x)=0$. Note $|\nabla u(z)|>0$, $|u(z)|<1$ and $|\nabla\sin^{-1}u(z)|^{2}=\lambda$; from Proposition 3.2, we get the extended splitting domain on $M^{n}$.
The main conclusion follows by the above and the induction on $\mathcal{N}(u)$.
Acknowledgments
We thank Xiaodong Wang for his comments on the early version of this paper.